OpenStudy (anonymous):

If Kenton walks for 60 minutes at the rate of 3 mph and then runs for 15 minutes at the rate of 8 mph, how many miles will he travel?

OpenStudy (anonymous):

5 mph

OpenStudy (anonymous):

5 miles

OpenStudy (mertsj):

3+1/4(8) 3+2=5

OpenStudy (anonymous):

8=4 + 2 miles + 60 minutes at 3mph= 3 miles Total five

OpenStudy (anonymous):

:) THXS